# Are formal and functional linguistic mechanisms dissociated in language models?

### v3

@@ 4.2 Functional Tasks
Indirect Object Identification (IOI) is a Situation Modeling / World Knowledge task providing inputs like “When Mary and John went to the store, John gave a bottle of milk to” and expecting the output “Mary”. Its corrupted version replaces the second instance of John with an unrelated name like Bob. Task performance is measured via the difference in the logit assigned to the correct vs. incorrect entity, i.e., ${logit}(\text{Mary})-{logit}(\text{John})$. This task, common throughout the circuits literature, requires models to recognize that, if an individual has an object, they cannot give it to themselves; this can be solved via situation modeling or potentially world knowledge. This task’s data is adapted from Wang et al. (2023), who introduced it and studied its circuit in GPT-2 small (Radford et al., 2019).
@@
Entity Tracking is a Situation Modeling task that gives models inputs like $s=$ “The apple is in Box F, the computer is in…the document is in Box Q. Box F contains the” and expects outputs like “apple” (Kim and Schuster, 2023). In the corrupted version, the queried object (e.g. computer in Box Q) is different. Task performance is measured via ${logit}(\text{apple}|s)-{logit}(\text{computer}|s)$. In this task, LLMs must track entities and their state over the length of a discourse. Prakash et al. (2024) studied this task in Llama-7B (Touvron et al., 2023).
@@
Colored Objects is a Situation Modeling task that gives models inputs like $s=$ “On the table, I see an orange textbook, a red puzzle, and a purple cup. What color is the textbook?”; the expected output is “orange”. In the corrupted version, object colors and the queried object (e.g. “blue mug”) are different. Task performance is measured via ${logit}(\text{orange}|s)-{logit}(\text{blue}|s)$. Merullo, Eickhoff, and Pavlick (2024) first studied this task in GPT-2 medium (Radford et al., 2019), showing that its circuit uses many of the same mechanisms as IOI does; however we adapt this task’s data from the original BigBench task (Srivastava et al., 2023).
@@
Greater-Than is a Formal Reasoning task with inputs like $s=$“The war lasted from the year 1842 to the year 18”; we expect outputs greater than 42. Performance is measured via $\sum_{y>\texttt{YY}}p(y|s)-\sum_{y\leq\texttt{YY}}p(y|s)$. The corrupted input for this task replaces the start year YY (in $s$, YY=“42”) with “01”, shifting the model’s output distribution towards years < YY. This task’s data is adapted from Hanna, Liu, and Variengien (2023), who introduced it, and studied its circuit in GPT-2 small (Radford et al., 2019).
@@
Country-Capital is a World Knowledge task that gives models inputs like $s=$ “France, whose capital, ” and expects outputs like “Paris”. Its corrupted variant replaces France with another country like Italy; task performance is then measured via ${logit}(\text{Paris}|s)-{logit}(\text{Rome}|s)$. We also include the reverse task (Capital-Country), which queries the country of which a city is the capital; past work has shown that this has a similar circuit to the Country-Capital task (Hanna, Pezzelle, and Belinkov, 2024).
@@ 5 Experimental Pipeline
Our experimental pipeline works as follows. For a given model, we find the circuit for each task in Section 4, using 500 examples per task. Then, for each pair of task circuits, we measure their similarity; if a formal-functional dissociation is to exist, formal circuits should be dissimilar from functional ones but similar to one another. Here, we discuss our choice of models, circuit-finding methods, and metrics to measure circuit similarity.
@@
We study state-of-the-art models from five families: Llama-3 8B (Llama Team, 2024), Gemma-2 2B (Gemma Team, 2024), Qwen-2 7B (Yang et al., 2024), Mistral-v0.3 7B (Jiang et al., 2023), and OLMo 7B (Groeneveld et al., 2024). We choose these models, lying in the 2-8 billion parameter range, because they are the largest models for which current circuit-finding techniques can function, due to both memory and compute constraints. Moreover, these models are capable enough to perform some situation modeling and world knowledge tasks, on which smaller models generally fail. Note that these are base models that underwent no instruction tuning, as this type of model was analyzed by Mahowald et al. and is commonly studied in the circuits literature. We speculate that results for instruction-tuned models would be similar, as recent work has shown that learned features are similar across base and instruction-tuned models (Kissane et al., 2024), and that fine-tuning mostly enhances existing circuits (Prakash et al., 2024).
@@
For each task, we estimate the IE of each edge in the model’s computational graph using EAP-IG, as described in Section 3.2. Each node in the computational graph represents an attention head, an MLP, and the model’s inputs or logits, and edges indicate causal links between nodes. Given these IEs, we can then construct a circuit by taking the top-$n$ edges with the highest absolute IE. For each task, we search for the minimum $n$ such that the top-$n$ circuit has a faithfulness of at least 85%, i.e., we find the minimum $n$ such that the circuit recovers at least 85% of the full model’s task performance. In this way, we find a circuit that explains almost all of the model’s performance on the task, without including a long tail of low-IE nodes and edges.
@@
Note that IoU and recall behave differently when circuits are of different sizes. If $|C_{1}|>>|C_{2}|$, IoU penalizes this heavily, as it is capped at $|C_{2}|/|C_{1}|$. While recall$(C_{2},C_{1})$ is also capped at $|C_{2}|/|C_{1}|$, recall$(C_{1},C_{2})$ can be as high as 1; indeed, large circuits may naturally recall more edges from smaller ones.
@@ 6 Do formal and functional networks overlap in LLMs?
In this section, we investigate the similarity of circuits for formal and functional linguistic tasks based on the degree to which the circuits’ edges overlap. In Section 6.1, we conduct an analysis of circuit overlap based on intersection over union. Then, we investigate whether functional task circuits may contain formal task circuits (Section 6.2). Finally, in Section 6.3 we characterize the edges that are shared between circuits.
@@ 6.1 Main experiment
In Figure 3, we report the IoU between each pair of task circuits, averaged across models. We also report the average percentage of the model’s edges included in each circuit, as the circuits vary in size: the SVA circuit contains only 0.71% of model edges, on average, while the Colored Objects task contains over 4x more, at 3.00%.
@@
Whether these IoUs are (statistically) significant depends on how we model a random, baseline circuit. Past work has considered random circuits with $n$ edges constructed by selecting them uniformly at random (Hanna, Pezzelle, and Belinkov, 2024; Shi et al., 2024). In this case, we can model the probability of two such circuits having an overlap of a given size using a hypergeometric distribution (App. B), and near all of the overlaps in Figure 3 are significantly higher than chance.
@@
Few tasks have even moderate overlap. Capital-Country and Country-Capital have high IoU (0.43), likely because they are nearly the same task; they share the same structure (fact-retrieval) and domain (geography). Similarly, IOI, Colored Objects, and Entity Tracking all have moderately high overlap (an average IoU of 0.27), perhaps because they all involve situation modeling; Merullo, Eickhoff, and Pavlick (2024) also found that the IOI and Colored Objects circuits overlap.
@@
We can verify these groupings found by visual inspection using clustering. We use agglomerative clustering to find which groups of tasks are closest to one another, as measured by the Euclidean distance of their IoU vectors (i.e. the rows or columns of Figure 3, left). We then create a dendrogram using the results of this clustering. The resulting dendrogram (Figure 3, right) shows that the Capital-Country / Country-Capital grouping and the IOI / Colored Objects / Entity Tracking groupings are also found by the clustering algorithm. It also finds a purely formal cluster that is less visible on the heatmap: Wug Test / SVA / NPI. However, the broader clusters do not reflect a formal-functional split. The formal Wug Test / SVA / NPI cluster next merges with Greater-Than (functional), and the functional IOI / Colored Object / Entity-Tracking cluster merges with Hypernymy (formal).
@@
The results of this experiment suggest that there is low but non-zero overlap between circuits for formal and functional linguistic abilities. At the same time, the formal tasks do not have especially high overlap with one another. Taken together, these two facts weigh against a potential formal-functional distinction. However, there remain open questions: could the overlap between formal and function task circuits stem from the fact that both categories of tasks are expressed in language? Moreover, what is the nature of the overlap between the formal and functional tasks? If that low overlap corresponds to a task-agnostic shared processing mechanism, akin to low-level visual systems in the brain, the formal-functional hypothesis would be more plausible.
@@ 6.2 Do functional task circuits contain formal task circuits?
One way to demonstrate that functional tasks do not contain formal tasks is to perform the same analysis as above but measure recall, a directional measure of how much a given circuit contains another one. The results of this analysis (Figure 4, left) show that functional tasks indeed do not subsume formal ones. While some formal tasks (Gendered Pronoun and SVA) with small circuits are well captured by most other task circuits, other, formal tasks like Hypernymy and NPI, are not so well captured. Moreover, Hypernymy has high coverage of many functional tasks, while the reverse is not true. In general, trends in recall seem dominated by the size of the circuits. Circuits that are large (like Hypernymy, IOI, Colored Objects, and Entity Tracking) have high recall with respect to other circuits, while being poorly covered by other circuits. Formal task circuits are not always smaller than functional task circuits: there are large formal task circuits (Hypernymy) and small functional task circuits (Country-Capital).
We also perform a clustering analysis on the reference recall vectors, using the same settings as in Section 6.1; that is, the clustering should find groups of tasks that are well recalled by similar tasks. This analysis (Figure 4, right) yields similarly negative results. Formal tasks cluster together, but imperfectly: the formal cluster includes Greater-Than, as well as Capital-Country and Country-Capital, more loosely.
@@
We can also investigate whether the small formal-functional overlap we observe is due to our posing functional tasks in language, by analyzing two tasks that are purely functional, without any language involved. The first is Nikankin et al.’s (2024) basic Math task, which consists of simple arithmetic problems involving addition, subtraction, and multiplication, such as 25 + 3 =. Corrupted examples share the same operator, but different operands. We measure performance via the difference in logit assigned to the clean and corrupted equations’ answers. Presented in purely symbolic form, this task should engage purely functional (formal reasoning) areas of the model, if such exist. This task is highly sensitive to tokenization, [^2] LLMs differ widely in how they tokenize numbers; while Llama-3 tokenizes numbers up to 3 digits in length as one token, other models (e.g. Gemma-2) tokenize them digit-by-digit, and others tokenize them in groups of 2. Single-token prediction tasks are most compatible with circuit analysis, so we use Llama-3. and cannot easily be ported to new models, so we study it in Llama-3 (8B) alone.
The second is a string manipulation task called Echo, inspired by a task of the same name from Hupkes et al. (2021). In this task, models see 4 distinct tokens randomly sampled from the model’s tokenizer, and must repeat the last token; a successfully solved example looks like “$t_{1}$ $t_{2}$ $t_{3}$ $t_{4}$ : $t_{4}$”. We give the model two solved examples (2-shot), and then have it finish an incomplete example. In corrupted examples, the final token ($t_{4}$) is replaced with another, distinct token. We measure performance via the logit assigned to the clean and corrupted examples’ final token. Since the tokens are randomly sampled, there should be no trace of linguistic structure; this task is most akin to a pattern-recognition formal reasoning task.
We find the circuits for these two tasks, and compare them to our existing task circuits using IoU, as done previously. Our results on these two tasks (Figure 5) are similar to those of other functional tasks. Neither Math nor Echo is particularly similar to any other task. Notably, although one might hypothesize that they are supported by the same mechanisms, Math and Greater-Than do not have especially high overlap. Echo also overlaps slightly more with Colored Objects, IOI, Entity Tracking, and Hypernymy than with other tasks. In general, IoU between Math / Echo and formal tasks is low but non-zero (0.09 on average). We take this as evidence that the overlap between formal and functional linguistic tasks is not due to the latter containing the former.

### v4

@@ 4.2 Functional Tasks
Indirect Object Identification (IOI) is a Situation Modeling / World Knowledge task providing inputs like “When Mary and John went to the store, John gave a bottle of milk to” and expecting the output “Mary”. Its corrupted version replaces the second instance of John with an unrelated name like Bob. Task performance is measured via the difference in the logit assigned to the correct vs. incorrect entity, i.e., ${logit}(\text{Mary})-{logit}(\text{John})$. This task, common throughout the circuits literature, requires models to recognize that, if an individual has an object, they cannot give it to themselves. This can be solved via situation modeling, which entails recognizing that Mary is the other entity in this situation, along with the commonsense knowledge that people seldom give things to themselves; in Mahowald et al.’s (2024) framework, this is considered world knowledge. In neuroscience, such commonsense world knowledge has been found to exist outside the language network (Ivanova et al., 2021). This task’s data is adapted from Wang et al. (2023), who introduced it and studied its circuit in GPT-2 small (Radford et al., 2019).
@@
Entity Tracking is a Situation Modeling task that gives models inputs like $s=$ “The apple is in Box F, the computer is in…the document is in Box Q. Box F contains the” and expects outputs like “apple” (Kim and Schuster, 2023). In the corrupted version, the queried object (e.g. computer in Box Q) is different. Task performance is measured via ${logit}(\text{apple}|s)-{logit}(\text{computer}|s)$. In this task, highlighted by Mahowald et al. as an example of situation modeling, LLMs must track entities and their state over the length of a discourse; past work in neuroscience has studied the requisite situation modeling skills (Baldassano et al., 2017). Prakash et al. (2024) studied this task in Llama-7B (Touvron et al., 2023).
@@
Colored Objects is a Situation Modeling task that gives models inputs like $s=$ “On the table, I see an orange textbook, a red puzzle, and a purple cup. What color is the textbook?”; the expected output is “orange”. In the corrupted version, object colors and the queried object (e.g. “blue mug”) are different. Task performance is measured via ${logit}(\text{orange}|s)-{logit}(\text{blue}|s)$. Much like the preceding task, the Colored Objects task requires the LLM to model the situation at hand and recall facts about it, making it a situation modeling task; past work has localized situation modeling abilities to the brain’s default network, outside the language network (Randy L. Bruckner, 2019). Merullo, Eickhoff, and Pavlick (2024) first studied this task in GPT-2 medium (Radford et al., 2019), showing that its circuit uses many of the same mechanisms as IOI does. We adapt this task’s data from the original BigBench task (BigBench Team, 2023).
@@
Greater-Than is a Formal Reasoning task with inputs like s=s=“The war lasted from the year 1842 to the year 18”; we expect outputs greater than 42. Performance is measured via $\sum_{y>\texttt{YY}}p(y|s)-\sum_{y\leq\texttt{YY}}p(y|s)$. The corrupted input for this task replaces the start year YY (in $s$, YY=“42”) with “01”. Corrupted inputs thus shift the model’s output distribution towards early years like “02” or “03”; these are in general $\leq$ YY, as the start year YY in clean inputs ranges from “02” to “98”. We adapt this task’s data from Hanna, Liu, and Variengien (2023), who introduced it, and studied its circuit in GPT-2 small. Performing the comparison involved in this task requires functional competence; for similar work in humans, see Amalric and Dehaene (2019), who study math presented both in symbols and in natural language.
@@
Country-Capital is a World Knowledge task that gives models inputs like $s=$ “France, whose capital,” and expects outputs like “Paris”. Its corrupted variant replaces France with another country like Italy; task performance is then measured via ${logit}(\text{Paris}|s)-{logit}(\text{Rome}|s)$. We also include the reverse task (Capital-Country), which provides a city, and asks which country it is the capital of; past work has shown that this has a similar circuit to the Country-Capital task (Hanna, Pezzelle, and Belinkov, 2024). Both of these tasks involve factual / world knowledge, handled by the brain’s knowledge and reasoning systems, exterior to the language network (Fedorenko, Ivanova, and Regev, 2024).
@@ 5 Experimental Pipeline
Our experimental pipeline works as follows. For a given model, we find the circuit for each task in Section˜4, using 500 examples per task. Then, for each pair of task circuits, we measure their similarity; if a formal-functional dissociation is to exist, formal circuits should be dissimilar from functional ones but similar to one another. Here, we discuss our choice of models, circuit-finding methods, and metrics to measure circuit similarity.
@@
We study state-of-the-art models from five families: Llama-3 8B (Llama Team, 2024), Gemma-2 2B (Gemma Team, 2024), Qwen-2 7B (Yang et al., 2024), Mistral-v0.3 7B (Jiang et al., 2023), and OLMo 7B (Groeneveld et al., 2024). We choose these models, lying in the 2-8 billion parameter range, because they are the largest models for which current circuit-finding techniques can function, due to both memory and compute constraints. Moreover, these models are capable enough to perform some situation modeling and world knowledge tasks, on which smaller models generally fail. Note that these are base models that underwent no instruction tuning, as this type of model was analyzed by Mahowald et al. and is commonly studied in the circuits literature. We speculate that results for instruction-tuned models would be similar, as recent work has shown that learned features are similar across base and instruction-tuned models (Kissane et al., 2024b), and that fine-tuning mostly enhances existing circuits (Prakash et al., 2024).
@@
For each task, we estimate the IE of each edge in the model’s computational graph using EAP-IG, as described in Section˜3.2. Each node in the computational graph represents an attention head, an MLP, and the model’s inputs or logits, and edges indicate causal links between nodes. Given these IEs, we can then construct a circuit by taking the top-$n$ edges with the highest absolute IE. For each task, we search for the minimum $n$ such that the top-$n$ circuit has a faithfulness of at least 85%, i.e., we find the minimum $n$ such that the circuit recovers at least 85% of the full model’s task performance. In this way, we find a circuit that explains almost all of the model’s performance on the task, without including a long tail of low-IE nodes and edges. [^3] See Appendix D for results with a higher threshold, 90%, yielding the same results as with the lower, 85% threshold.
@@
[…]
@@
One potential concern in designing overlap-based metrics for circuit similarity is superposition, along with polysemanticity. Superposition is typically discussed as the phenomenon in which LLMs represent more features than they have dimensions (or neurons) (Elhage et al., 2022). Since the number of features greater than the number of neurons, some neurons must become polysemantic: they fire on many different features in the input, rather than always firing on or representing just one feature. Later work has observed that this phenomenon is not constrained to neurons: attention heads, too, can be polysemantic (Kissane et al., 2024a). This raises questions, namely: if two task circuits overlapped, but this overlap were due only to polysemanticity, would our methods catch this? And does this matter?
It is possible that two task circuits could overlap due to polysemanticity, and our methods take no explicit steps to exclude overlaps of this type. Consider, though, that the ideas underlying superposition come from neuroscience: past work has hypothesized that the human brain also represents more features than it has neurons (Olshausen and Field, 1997). Furthermore, the localizer methodology used to find the language network also contains no safeguards against overlap due to superposition. Despite this, there is no consistent overlap between the language network and other networks across tasks and individuals. For this reason, we hold LLMs to a similar standard, requiring there to be no overlap between formal and functional circuits, irrespective of superposition.
@@ 6 Do formal and functional networks overlap in LLMs?
In this section, we investigate the similarity of circuits for formal and functional linguistic tasks based on the degree to which the circuits’ edges overlap. In Section˜6.1, we conduct an analysis of circuit overlap based on intersection over union. Then, we investigate whether functional task circuits may contain formal task circuits (Section˜6.2). Finally, in Section˜6.3 we characterize the edges that are shared between circuits.
@@ 6.1 Main experiment
In Figure˜3, we report the IoU between each pair of task circuits, averaged across models. We also report the average percentage of the model’s edges included in each circuit, as the circuits vary in size: the SVA circuit contains only 0.71% of model edges, on average, while the Colored Objects task contains over 4x more, at 3.00%.
@@
Whether these IoUs are (statistically) significant depends on how we model a random, baseline circuit. Past work has considered random circuits with $n$ edges constructed by selecting them uniformly at random (Hanna, Pezzelle, and Belinkov, 2024; Shi et al., 2024). In this case, we can model the probability of two such circuits having an overlap of a given size using a hypergeometric distribution, and near all of the overlaps in Figure˜3 are significantly higher than chance; see Appendix B for details on this and other techniques for modeling random overlap, all of which judge these IoUs to be significant.
@@
Few tasks have even moderate overlap. Capital-Country and Country-Capital have high IoU (0.43), likely because they are nearly the same task; they share the same structure (fact-retrieval) and domain (geography). Similarly, IOI, Colored Objects, and Entity Tracking all have moderately high overlap (an average IoU of 0.27), perhaps because they all involve situation modeling; Merullo, Eickhoff, and Pavlick (2024) also found that the IOI and Colored Objects circuits overlap. Note that even these moderate values are relatively high for edge IoU, as rather similar task circuits may only have up to 0.3–0.4 IoU; see Appendix C for more details on the ranges of edge IoU values for similar circuits.
@@
We can verify these groupings found by visual inspection using clustering. We use agglomerative clustering to find which groups of tasks are closest to one another, as measured by the Euclidean distance of their IoU vectors (i.e. the rows or columns of Figure˜3, left). We then create a dendrogram using the results of this clustering. The resulting dendrogram (Figure˜3, right) shows that the Capital-Country / Country-Capital grouping and the IOI / Colored Objects / Entity Tracking groupings are also found by the clustering algorithm. It also finds a purely formal cluster that is less visible on the heatmap: Wug Test / SVA / NPI. However, the broader clusters do not reflect a formal-functional split. The formal Wug Test / SVA / NPI cluster next merges with Greater-Than (functional), and the functional IOI / Colored Object / Entity-Tracking cluster merges with Hypernymy (formal).
It is notable that, of all the formal tasks, hypernymy is the one that is misgrouped. This is in some sense consonant with its controversially formal status. While Mahowald et al. and similar researchers might categorize such lexical semantic relations, this is not a universal view. Others find less of a distinction between lexical semantics and, e.g., world-knowledge semantics (Hagoort et al., 2004), or even view hypernymy as belonging to the latter group, more related to reasoning than lexical semantics (Benn et al., 2023). From these perspectives, the misgrouping of hypernymy could stem from the fact that it is not a formal linguistic task at all, but a functional one. However, we note that this account leaves unexplained the fact that Greater-Than, an uncontroversially functional task, is grouped with the formal tasks.
@@
Ultimately, the results of this experiment suggest that there is low but non-zero overlap between circuits for formal and functional linguistic abilities. At the same time, the formal tasks do not have especially high overlap with one another. Taken together, these two facts weigh against a potential formal-functional distinction. However, there remain open questions: could the overlap between formal and function task circuits stem from the fact that both categories of tasks are expressed in language? Moreover, what is the nature of the overlap between the formal and functional tasks? If that low overlap corresponds to a task-agnostic shared processing mechanism, akin to low-level visual systems in the brain, the formal-functional hypothesis would be more plausible.
@@ 6.2 Do functional task circuits contain formal task circuits?
One way to demonstrate that functional tasks do not contain formal tasks is to perform the same analysis as above but measure recall, a directional measure of how much a given circuit contains another one. The results of this analysis (Figure˜4, left) show that functional tasks indeed do not subsume formal ones. While some formal tasks (Gendered Pronoun and SVA) with small circuits are well captured by most other task circuits, other, formal tasks like Hypernymy and NPI, are not so well captured. Moreover, Hypernymy has high coverage of many functional tasks, while the reverse is not true. In general, trends in recall seem dominated by the size of the circuits. Circuits that are large (like Hypernymy, IOI, Colored Objects, and Entity Tracking) have high recall with respect to other circuits, while being poorly covered by other circuits. Formal task circuits are not always smaller than functional task circuits: there are large formal task circuits (Hypernymy) and small functional task circuits (Country-Capital).
We also perform a clustering analysis on the reference recall vectors, using the same settings as in Section˜6.1; that is, the clustering should find groups of tasks that are well recalled by similar tasks. This analysis (Figure˜4, right) yields similarly negative results. Formal tasks cluster together, but imperfectly: the formal cluster includes Greater-Than, as well as Capital-Country and Country-Capital, more loosely.
@@
We can also test whether the small formal-functional overlap we observe is due to our posing functional tasks in language, by analyzing versions of our functional tasks that are purely functional, without any language involved. We thus create non-language-mediated versions of every one of our functional tasks; we call these our purely-functional tasks:
  • IOI (Situation Modeling): we alter the task to give models inputs like $s=$ “Alice Bob Alice Bob. Carrie Dylan Dylan Carrie. Mary John John” and expect outputs like “Mary”. That is, the model must repeat the name that has not already been repeated, as in the original IOI task. Note that models are given in-context examples, to allow them to understand the task and format. Unlike the following purely functional tasks, this purely- functional version of IOI diverges relatively strongly from the original, which lies at the intersection of various functional capabilities such as Situation Modeling and World Knowledge; the purely-functional version does not test the latter.
  • Entity Tracking (Situation Modeling): we alter the task to give models inputs like s=s=apple : G, computer : F, rock : A; F : computer. ring : A, note : B, book : P, glass : W; A :” and expect outputs like “ring”. That is, the task consists of a list of comma-separated “object : letter” pairs, followed by a semicolon and letter; models must report the object paired with the letter in the list. Models are given one in-context example (shown in italics), to help them understand the task format.
  • Colored Objects (Situation Modeling): we alter the task to give models inputs like $s=$ “orange textbook, red puzzle, purple cup; textbook:”; the expected output is “orange”. Models are given one in-context example (not shown), to help them understand the task format. Note that in this distilled version, Colored Objects is near identical to Entity Tracking (though not IOI), reflecting the similarity of the two original tasks.
  • Greater-Than (Formal Reasoning): we alter the task to give models inputs like s=s=“1842 < 18”; we expect outputs greater than 42.
  • Country-Capital (World Knowledge): we alter the task to give models inputs like $s=$ “The Netherlands : Amsterdam :: France :”; we expect outputs like “Paris”. We include an in-context example demonstrating the task format. We also include the reverse task (Capital-Country), using the same analogy structure, with the relation reversed.
We study all of these tasks only in Llama-3 (8B), the most capable of our models; these abstract, in-context tasks are somewhat challenging for smaller models. We find the circuits for these five tasks, and compare them to our existing task circuits using IoU, as done previously. Here, we compare formal and purely-functional tasks, to test whether the (small) overlap we observed previously was due to using stimuli in the language format. If that was the case, that overlap should disappear in this more controlled setup. See Appendix D for comparison of the original functional tasks and their purely-functional counterparts. For more purely-functional tasks, see Appendix B.
Figure 5: Edge IoU (left) and recall (right) heatmaps between formal and purely function (PF) tasks. Notably, the use of purely functional tasks has changed little, aside from weaker similarity between IOI and Colored Objects / Entity Tracking.
Our results on these five purely-functional tasks (Figure˜5) are similar to those obtained using our original functional tasks. The same groupings appear: Country-Capital and Capital-Country are similar, and Colored Objects and Entity Tracking are as well, though IOI is now less similar to them. In general, IoU between purely-functional and formal tasks is still low but non-zero. We take this as evidence that the overlap between formal and functional linguistic tasks is not due to the latter containing the former.